OpenStudy (anonymous):

If the x and y are the tens digits and the units digit resp.,of the product 725,278 x 67066,what is the value of x+y?

OpenStudy (anonymous):

multiply 752,278 x 67,066 48,641,494,348 is the answer the ones digit, y, is 8 the tens digit, x, is 4 x+y = 4+8 =12

OpenStudy (asnaseer):

since you are only interested in the last two digits, you only need to multiply the last two digits of each number:\[78*66=5148\]so x=4, y=8